OpenStudy (anonymous):

You have a total of 21 coins, all nickles and dimes. The total value is $1.70. Write and solve a system of equations to find the number of dimes d and the number of nickles n that you have. Please help me step by step.

OpenStudy (anonymous):

d(10)+n(5)=1.50

OpenStudy (anonymous):

solve for d and n

OpenStudy (anonymous):

you mean 1.70

OpenStudy (anonymous):

Oh, yes. Sorry.

OpenStudy (anonymous):

its ok sooo... it would look like this

OpenStudy (anonymous):

d + n=1.70 d(10) + n(5)=170

OpenStudy (anonymous):

I think it would be d+n=21 d(10)+n(5)=1.70

OpenStudy (anonymous):

yeah okay i forgot about the 21

OpenStudy (anonymous):

Hold on im going to work it out... then ask u if its right.

OpenStudy (anonymous):

okie dokie

OpenStudy (anonymous):

ok i did something wrong.... i got 20.66 for d

OpenStudy (anonymous):



OpenStudy (mathstudent55):

You have a small mistake that changes the answer completely.

OpenStudy (mathstudent55):